# Smoothing of Headland Path Edges and Headland-to-Mainfield Lane Transitions Based on a Spatial Domain Transformation and Linear Programming

### v3

@@ 5 Discussion
A brief comment about path planning methods by neural networks is made (see a simplistic algorithm in Graf Plessen (2020) and the references therein). The key argument against the employment of neural networks is that both path planning Problems 1 and 2 deal with an infinite number of possible tasks for headland path edges and headland-to-mainfield lane transitions and require a continuous action space. To encode a path planning module in a neural network an enormous number of training examples and their ground-truth solutions would be required, since there is an enormous variety in nature of different field contours, which would make a large training set mandatory. To underline the continous control aspect, exemplarily see Plappert et al. (2018) where policies for continuous control are learned. However, each of the evaluation problems receives a reward either (i) upon moving forward (HalfCheetah, Hopper, Swimmer, Walker2D, SparseHalfCheetah, SwimmerGather), or (ii) upon reaching a quasi-stationary goal position ( InvertedDoublePendulum, InvertedPendulum, Reacher, SparseCartpoleSwingup, SparseMountainCar, SparseDoublePendulum). There is no option for path shaping by accounting for precision constraints. Training time and hardware requirements are not stated. In contrast, proposed method is based on linear programming, with hyperparameter-free objective function, fast solve times, and can incorporate multivariate nonlinear vehicle dynamics and constraints on both states and control inputs in a structured manner to allow flexibility and generate deterministic precision paths, which is required for both Problems 1 and 2.
@@
Peng et al. (2023) propose an optimisation‐based algorithm in the time domain for headland turning under constraints imposed by headland geometry and obstacles, especially in orchard-like environments. In our paper obstacles in the headland are not assumed, however, since being a relevant problem the topic of obstacle avoidance within the spatial domain is briefly addressed. Based on previous work it is pointed to three options: (i) using a geometric path planner to solve the combinatorial corridor planning problem (Plessen et al. (2017a)), (ii) using a linearisation-approach to model obstacles in the spatial domain (Plessen et al. (2017b)), and (iii) using path corridor heuristics (Plessen (2017)). The last reference further discusses how velocity bounds within friction limits could be incorporated into spatial-based optimisation problems.
@@
Finally, a comment about manual instead of autonomous driving is made. A visualisation such as Fig. 8 gives a farmer a reference path plan and practical guidance for how to steer. The importance of generated path plans for manual driving should not be underestimated, since the decision about how to drive transitions between headland path and mainfield lanes is often not straightforward in practice.
@@ 6 Conclusion
This paper contributed to the task of path planning within agricultural fields by proposing a method based on a spatial domain transformation of vehicle dynamics and linear programming to address the two tasks of smoothing of headland path edges and headland-to-mainfield lane transitions. Within the agricultural context this is the first paper suggesting to use a spatial domain vehicle dynamics representation in combination with a numerical optimisation approach to compute precision paths. Proposed method is of interest for both autonomous robot applications as well as as to provide high-level reference path plans for manual driving.
This paper is limited to forward-motion only. Methods can be extended to also permit backward-motion in the path planning, e.g., to permit three-point steering maneuvers involving both forward and backward motion at the transitions between headland path and mainfield lanes or headland path edges.
The main avenue of future work is the development of alternative heuristics to the 5-point piecewise-affine reference generation step for Problem Type 1.

### v4

@@ 6 Discussion
A brief comment about path planning methods by neural networks is made (see a simplistic algorithm in Plessen (2020) and the references therein). The key argument against the employment of neural networks is that both path planning Problems 1 and 2 deal with an infinite number of possible tasks for headland path edges and headland-to-mainfield lane transitions and require a continuous action space. To encode a path planning module in a neural network, an enormous number of training examples and their ground-truth solutions would be required, since there is an enormous variety in nature of different field contours. This would make a large training set mandatory. To underline the continous control aspect, see Plappert et al. (2018) where policies for continuous control are learned. However, each of the evaluation problems receives a reward either (i) upon moving forward (HalfCheetah, Hopper, Swimmer, Walker2D, SparseHalfCheetah, SwimmerGather), or (ii) upon reaching a quasi-stationary goal position
(InvertedDoublePendulum, InvertedPendulum, Reacher,
SparseCartpoleSwingup, SparseMountainCar, SparseDoublePendulum).
There is no option for path shaping by accounting for precision constraints. Training time and hardware requirements are not stated. In contrast, the proposed method here is based on linear programming, with a hyperparameter-free objective function, fast solution times, and can incorporate multivariate nonlinear vehicle dynamics and constraints on both states and control inputs in a structured manner to allow flexibility. Therefore, the proposed method can generate deterministic precision paths, which are required for both Problems 1 and 2.
@@
Peng et al. (2023) proposed an optimisation‐based algorithm in the time domain for headland turning under constraints imposed by headland geometry and obstacles, especially in orchard-like environments. In this paper, obstacles in the headland are not assumed, however, since obstacles are a relevant problem, the topic of obstacle avoidance within the spatial domain is briefly addressed. Based on previous work, three options have been proposed: (i) using a geometric path planner to solve the combinatorial corridor planning problem (Plessen et al. (2017a)), (ii) using a linearisation-approach to model obstacles in the spatial domain (Plessen et al. (2017b)), and (iii) using path corridor heuristics (Plessen (2017)). The last reference further discusses how velocity bounds within friction limits could be incorporated into spatial-based optimisation problems.
@@
Finally, a comment about manual instead of autonomous driving is made. A visualisation, such as Fig. 8, gives a farmer a reference path plan and practical guidance for how to steer. The importance of generated path plans for manual driving should not be underestimated, since the decision about how to drive transitions between a headland path and mainfield lanes is often not straightforward in practice.
@@ 7 Conclusion
This paper contributed to the task of path planning within agricultural fields by proposing a method based on a spatial domain transformation of vehicle dynamics and linear programming to address the two tasks of smoothing of headland path edges and headland-to-mainfield lane transitions. Within the agricultural context this is the first known approach suggesting to use a spatial domain vehicle dynamics representation in combination with a numerical optimisation approach to compute precision paths. The proposed method is of interest for both autonomous robot applications as well as as to provide high-level reference path plans for manual driving.
The main avenue of future work is the development of alternative heuristics to the 5-point piecewise-affine reference generation step for Problem Type 1. A second topic is the extension of the methods to also permit backward-motion in the path planning. This would permit three-point steering maneuvers involving both forward and backward motion at the transitions between headland path and mainfield lanes or headland path edges.